# Detecting Throat Cancer from Speech Signals using Machine Learning: A Scoping Literature Review

### v2

@@ 3.3 RQ2 - Feature Extraction
We present the features extracted from the examined articles and discuss how they perform. In Figure 7, we show the number of each feature extraction method used in the articles. For clarity, methods used only once are grouped into the ’other’ category. Some articles combined features as input into classification systems although this is not shown in Figure 7.
@@
Figure 8 shows the results obtained from methods using different features. Figure 8(a) shows a disparity between the two most common feature sets in binary classification, with acoustic features performing consistently worse than MFCC. Glottal parameters seem to generally perform the best, although their performance still lies within the boundary of the articles using MFCC. As there is only one example of raw signals and STFT being used, it is hard to comment on their performance compared to the other methods.
Figure 8(b) shows the UAR achieved using different features in the multi-class tasks. These notably use a less wide range of features than the binary classification tasks, with a combination of features being used in many of the articles. Combining features seems to perform slightly better than MFCCs alone, although the performances are not significantly different.
As stated in the previous section, comparing the results of articles that use different datasets is hard. However, we can compare the results of the articles submitted to the IEEE FEMH Voice Data Challenge 2018. Figure 9 shows the results obtained based on the features used. In the challenge articles, only two types of features were used: MFCC and a combination of different features. This combination generally included MFCC, acoustic features, and spectral features. There is no clear advantage to either of these feature inputs.
From these results, we can see that no single feature set performs the best. For binary classification, glottal parameters have the best average accuracy, although this still lies within the range of results found when using MFCCs. The results of the IEEE FEMH Voice Data Challenge 2018 does not show any advantage when combining features over using MFCC alone.
@@
(b) The UAR achieved in the multi-class classifiers using different features
@@
Figure 9: The UAR found in the articles submitted to the IEEE FEMH Voice Data Challenge 2018 for different input features [44, 46, 45, 40, 48, 42, 41, 49, 47, 43].
@@ 3.4 RQ3 - Strengths, Weaknesses and Recomendations
Kim et al. [16] 50% 71% 1% 60% 40%
@@
Godino-Llorente and Gomez-Vilda [33] 25% 44% 0% 40% 0%
Ezzine et al. [35] 50% 29% 1% 20% 20%
Fang et al. [36] 50% 44% 0% 40% 60%
Ben Aicha and Ezzine [34] 25% 17% 0% 0% 20%
Ritchings et al. [32] 0% 12% 1% 0% 0%
Gavidia-Ceballos and Hansen [31] 25% 28% 0% 0% 20%
Kwon et al. [18] 100% 67% 1% 20% 60%
Grzywalski et al. [44] 50% 47% 0% 20% 0%
Miliaresi et al. [51] 25% 41% 1% 20% 40%
Fang et al. [50] 25% 50% 1% 60% 80%
Verikas et al. [39] 50% 39% 2% 40% 40%
Ju et al. [46] 50% 44% 0% 20% 40%
Islam et al. [45] 75% 44% 1% 20% 0%
Arias-Londoño et al. [40] 75% 39% 1% 20% 40%
Pham et al. [48] 75% 39% 2% 20% 40%
Chuang et al. [42] 50% 33% 2% 20% 20%
Bhat and Kopparapu [41] 25% 28% 0% 20% 0%
Ramalingam et al. [49] 0% 22% 2% 0% 20%
Pishgar et al. [47] 75% 33% 3% 20% 20%
Degila et al. [43] 75% 33% 0% 20% 60%
Wang et al. [38] 50% 50% 11% 20% 80%
Kim et al. [28] 75% 72% 9% 60% 60%
Chen et al. [29] 25% 56% 6% 60% 20%
Za’im et al. [37] 75% 72% 18% 60% 60%
@@ 3.4.1 Key Weaknesses
When comparing the papers to TRIPOD-AI we observed notable discrepancies between journal papers (average coverage of 18 points) and conference papers (average coverage of 11 points). We speculate that this disparity may arise from the typically stricter page limits imposed on conference papers compared to journal articles.
A critical area of weakness identified through TRIPOD-AI assessment is the lack of adherence to open science (points 18a-18f, and 19). Shockingly, only one of the rewieved articles has made its code available to the general public [30]; the lack of code sharing severely reduces the ability of other researchers to reproduce their results or validate their methods using external data. Furthermore, only three articles utilize open-source datasets [35, 38, 30]. Additionally, many articles fail to report conflicts of interest (point 20) and provide details on study protocol (point 22) or registration (point 22). While some articles mention data availability upon request, this practice is considered inadequate in promoting transparency.
@@
Another area where papers fall short when compared to the TRIPOD-AI checklist is the lack of discussion on the practical application of AI systems in real-world medical settings (points 27a-27c). While half of the articles discuss steps for further research, only one addresses the impact of poor-quality input data, and none of the articles discuss the user inputs to a possible system. This gap in research indicates a need for more comprehensive studies that consider the practical implementation of AI systems in real-world medical settings, although this may be due to the relative newness of this area of study.
Additionally, the dependence on small datasets is concerning. As stated in Section 3.1, the largest reported dataset is that of Ezzine et al. [35], although we believe that this is an outlier due to omitted preprocessing steps, meaning that this large dataset does not represent the number of participants. The next largest dataset is that of Kwon et al. [18], containing 491 patients. Excluding Ezzine et al. [35] and Ben Aicha and Ezzine [34] due to unclear dataset size, the mean size of the datasets across the articles is 207 participants. We also note that only two of the included articles performs any external validation of their models [36, 38]; combining this with the small sample size puts them at high risk of bias and overfitting [58].
@@ 4 Conclusion
Neural networks emerged as the most commonly employed algorithm, although other algorithms demonstrated comparable success in classifying patients. The most common features used were MFCCs, although a wide variety of features were explored without a clear superiority observed among them. The results found in these works show that this is a promising area and that the use of AI in the detection of speech pathologies should be further investigated. We discuss the weaknesses in this area, particularly the lack of open science and the lack of publicly available code and datasets. The absence of publicly available code or datasets prevents the reproducibility and extension of findings across these articles. We also make recommendations for future research in this area to promote better open science and clarity of reporting. In future work, we will address this shortcoming, with the aim to create public code repositories, hence enabling external validation and facilitating improvements in subsequent studies.

### v3

@@ 3.3 RQ2 - Feature Extraction
We present the features extracted from the examined articles and discuss how they perform. An average of 1.5 input features were used in each article with Kim et al. [16] using the most input features (raw signal, MFCC, acoustic features, spectral features).
In Figure 9, we show the number of each feature extraction method used in the articles. For clarity, methods used only once are grouped into the ’other’ category. Some articles combined features as input into classification systems although this is not shown in Figure 9.
@@
Figure 10 shows the results obtained from methods using different features. Figure 10(a) shows the results obtained for binary classification. MFCC has the highest average performance, with the raw signal having the worst average performance. The poor performance of using a raw signal as input may be due to the large dimensionality of an audio signal, without extracting features, there will be lots of noise and redundant data input into the system. Only two articles use raw signals as input to the system [16, 37]. Glottal parameters have the broadest range of results, although these results come from only two articles with shared authors [31, 32].
Figure 10(b) shows the results obtained for multi-class classification. Cosine transform coefficients have the best average performance although there is only a single instance of these being used as system input. Mel energies have the lowest average and lowest single performance, although a single article obtains these results [34]. The single best result is obtained by Song et al. [51] using a combination of features, this combination contains MFCC and acoustic features, as well as demographic features.
@@
(b) The accuracy achieved in the multi-class classifiers using different features
@@
As stated in the previous section, comparing the results of articles that use different datasets is hard. However, we can compare the results of the articles submitted to the IEEE FEMH Voice Data Challenge 2018. Figure 11 shows the results obtained based on the features used, this is reported as UAR as this was the metric most widely reported in the challenge articles. In the challenge articles, only two types of features were used: MFCC and a combination of different features. This combination generally included MFCC, acoustic features, and spectral features. Figure 11(a) shows the results of articles on cross-validation.
(a) The UAR achieved in the IEEE FEMH Voice Data Challenge 2018 evaluated on cross-validation using different system inputs
(b) The UAR achieved in the IEEE FEMH Voice Data Challenge 2018 evaluated on a holdout test set using different system inputs
Figure 11: The UAR found in the articles submitted to the IEEE FEMH Voice Data Challenge 2018 for different system inputs [43, 45, 44, 39, 47, 41, 40, 48, 46, 42].
@@ 3.4 RQ3 - Strengths, Weaknesses and Recomendations
Gavidia-Ceballos and Hansen [29] 25% 28% 0% 0% 20%
Godino-Llorente and Gomez-Vilda [30] 25% 44% 0% 40% 0%
Verikas et al. [38] 50% 39% 29% 40% 40%
Ben Aicha and Ezzine [31] 25% 17% 0% 0% 20%
Ezzine et al. [32] 50% 29% 14% 20% 20%
Arias-Londoño et al. [39] 75% 39% 14% 20% 40%
Bhat and Kopparapu [40] 25% 28% 0% 20% 0%
Chuang et al. [41] 50% 33% 14% 20% 20%
Degila et al. [42] 75% 33% 0% 20% 60%
Grzywalski et al. [43] 50% 47% 0% 20% 0%
Islam et al. [44] 75% 44% 14% 20% 0%
Ju et al. [45] 50% 44% 0% 20% 40%
Pham et al. [47] 75% 39% 14% 20% 40%
Pishgar et al. [46] 75% 33% 14% 20% 20%
Ramalingam et al. [48] 0% 22% 14% 0% 20%
Fang et al. [49] 50% 44% 0% 40% 60%
Fang et al. [33] 25% 50% 14% 60% 80%
Kim et al. [16] 50% 71% 29% 60% 40%
Miliaresi et al. [50] 25% 41% 14% 20% 40%
Kwon et al. [18] 100% 67% 14% 20% 60%
@@
Chen et al. [34] 25% 56% 14% 60% 20%
@@ 3.4.1 Key Weaknesses
When comparing the papers to TRIPOD-AI we observed notable discrepancies between journal papers (average coverage of 18 points) and conference papers (average coverage of 12 points). We speculate that this disparity may arise from the typically stricter page limits imposed on conference papers compared to journal articles.
A critical area of weakness identified through TRIPOD-AI assessment is the lack of adherence to open science (points 18a-18f, and 19). Shockingly, only one of the reviewed articles has made its code available to the general public [28]; the lack of code sharing severely reduces the ability of other researchers to reproduce their results or validate their methods using external data. Furthermore, only three articles utilize open-source datasets [28, 32, 37]. Additionally, many articles fail to report conflicts of interest (point 20) and provide details on study protocol (point 22) or registration (point 22). While some articles mention data availability upon request, this practice is considered inadequate in promoting transparency [59].
@@
Another area where articles fall short compared to the TRIPOD-AI checklist is the lack of discussion on the practical application of AI systems in real-world medical settings (points 27a-27c). While half of the articles discuss steps for further research, only one addresses the impact of poor-quality input data, and none of the articles discuss the user inputs to a possible system. This gap in research indicates a need for more comprehensive studies that consider the practical implementation of AI systems in real-world medical settings, although this may be due to the relative newness of this area of study.
Additionally, the dependence on small datasets is concerning. As stated in Section 3.1, the largest reported dataset is that of Ezzine et al. [32], although we believe that this is an outlier due to omitted preprocessing steps, meaning that this large dataset does not represent the number of participants. The next largest dataset is that of Wang et al. [37], containing 2000 patients. Excluding Ezzine et al. [32] and Ben Aicha and Ezzine [31] due to unclear dataset size, the mean size of the datasets across the articles is 306 participants. We also note that only three of the included articles perform any external validation of their models [18, 33, 37]; combining this with the small sample size puts them at high risk of bias and overfitting [60].
@@ 4 Conclusion
Neural networks emerged as the most commonly employed classification algorithm, although other algorithms demonstrated comparable success in classifying patients. The most common features used were MFCCs, although a wide variety of features were explored without a clear superiority observed among them. The results found in these works show that this is a promising area and that the use of AI in the detection of speech pathologies should be further investigated. We discuss the weaknesses in this area, particularly the lack of open science and the lack of publicly available code and datasets. The absence of publicly available code or datasets prevents the reproducibility and extension of findings across these articles. We also make recommendations for future research in this area to promote better open science and clarity of reporting. In future work, we will address this shortcoming, with the aim to create public code repositories, hence enabling external validation and facilitating improvements in subsequent studies.